media%2F194%2F194d8b14-ae95-47f9-9f7a-22
Celebrity Xpedition Siver Shadow siver wind SeaDream Wind Star Wind Sur wind Spird 6.4 2.1 a. Determine an estimated regression equation that can be used to predict the overall score gwen the score for Shore Ecursons (to 3 deoimals) Overall th. Consider the adation of the independent varstble Foodiining Develop the esmated regression eaton tc b st deamals), Overall Shore Exeursions + Food/Dinieg with a Shore Excursions score of 80 and a Food/Dg Sooe of 90 (to 2 decimais) 3/23/2019 0 Type here to search DOLL st
Answer #1

Overall Shore Excursion Food/Driving SUMMARY OUTPUT 94.2 93.2 92.8 91.2 90.7 90.4 90.2 90.1 89.3 89 89.2 88.9 88.8 87 87.4 86a. Overall=69.633+0.231 Shore Excursion

b.

2 Overall Shore Excursion Food/Driving SUMMARY OUTPUT 94.2 93.2 92.8 91.2 90.7 90.4 90.2 90.1 89.3 89 89.2 88.9 88.8 87 87.4Overall=44.780+0.250 Shore Excursion+0.255 Food/Driving

c. If Shore Excursion score=80, Food/Driving score=90 then Estimated overall score=44.780+0.250 *80+0.255*90=87.73